OpenStudy (anonymous):

-9 = 5/17n step by step thanks!!

OpenStudy (cwrw238):

ok multiply both sides by 17n: -9*17n = 5 divide both sides by - 9*17: n = 5 /(-9*17) - which you can work out using a calculator

OpenStudy (anonymous):

sweet thanks!!

OpenStudy (cwrw238):

yw